# The Whole Is Bigger Than the Sum of Its Parts: Modeling Individual Annotators to Capture Emotional Variability

### v1

@@ 2 Related Work
Instead, we build on the label processing method developed in previous work by Zhang et al. [7], which incorporated inter-annotator variance into machine learning models by creating new ground truth labels that incorporate this knowledge [7]. They upsampled existing annotations by selecting random subsets of annotators for each utterance and took the mean across those annotations. They added random noise to the resulting means, such that $x\ noise\sim U(-\frac{std(x)}{2},\frac{std(x)}{2})$ [^1] We also add $\epsilon=1E^{-12}$ to this value to account for cases where standard deviation is 0., where $x$ is activation or valence, $std$ indicates the standard deviation of the annotator ratings for that utterance, and $U$ is the uniform distribution. Kernel Density Estimation (KDE) via Diffusion was then calculated over the upsampled observations. They divided the KDE output grid into $N$ bins for each dimension and took the mean over the KDE samples inside each bin. They converted this grid to a probability distribution by normalizing over the means. The authors investigated $N=2$ and $N=4$. The KDE step was essential to remove sensitivity to where boundaries were drawn. The authors then trained a model to predict these binned distributions. However, in this approach, the model loses information about individual annotators. Additionally, because the approach is not differentiable, it cannot be included in model training. We present an approach with a differentiable component that permits learning a binned distribution, implemented using sigmoid-based soft operations.
@@ 3.1 Data setup
Dataset Preprocessing We process all datasets in the same way. We use min-max scaling on the annotator and consensus labels for activation and valence to restrict labels to the $[-1,1]$ range. We then use KDE to generate a 2D ground truth probability distributions as in [7]. We use a KDE grid size of 512 as we assume this will be sufficiently large to ensure the probability is insensitive to the grid boundaries.
@@ 3.2 Model Architecture
The baseline model directly learns the generated KDE distribution (as in [7]), having a final linear layer output of 16 logits for the 4x4 discretized KDE distribution prediction case. The MT model has separate prediction layers for each annotator as in [17]. Each annotator’s continuous prediction of activation and valence is made via a linear layer with an output size of 1. We use the one-hot method, previously used for text emotion recognition [20]. We use the same architecture as in the MT case but with only one annotator prediction head. The annotator ID is one-hot encoded and concatenated to the Wav2Vec2 and BERT embeddings on the model input. When training within corpus, we reduce computation cost by making predictions only for annotators in the batch input to the model.
@@ 3.3 Training Tasks
In this section, we define three different training tasks. The Baseline model is trained with the Baseline task. MT and one-hot models are trained by interleaving Tasks 1 and 2 (Task 1+2), defined below. We use stochastic gradient descent with a learning rate of $0.001$, with a learning rate scheduler that adjusts the learning rate by a factor of $0.1$ after five epochs of no reduction in validation metrics. We train models until early stopping triggers with a patience of 10 with a minimum of 30 epochs. Each model trains with a batch size of 32. For all methods we use the relevant task’s validation losses.
@@
Task 1 - Annotator Training: We train annotator-specific predictions using the individual annotator ground truth. We use Lin’s Concordance Correlation Coefficient (CCC) loss as our loss function since it better models dimensional attributes than other regression losses [21]. The sets $act$, and $val$ contain the ground truth labels from all annotators in a training batch. The sets $m_{act}$ and $m_{val}$ contain the model’s estimates of these labels. The loss is $2-CCC(m_{act},act)-CCC(m_{val},val)$.
Task 2 - DiffKDE: We learn the probability distribution using the KDE-generated ground truth labels. The model must produce a probability distribution from the model’s activation and valence predictions. However, the KDE method outlined by Zhang et al. in [7] is not immediately usable. KDE via diffusion starts with a histogram [22]. For each annotation we must know if it is in a particular bin to increment the bin’s histogram count. This operation is a binary operation and not differentiable. We introduce a differentiable approximation to this problem by instead calculating a confidence value that a given annotation is within a given bin. We modify an existing one-dimensional (1D) soft-histogram [^3] https://discuss.pytorch.org/t/differentiable-torch-histc/25865/4, as below, for the 2D data.
We use 64 bins for DiffKDE [^4] Note: smaller than 512 (used to generate target labels) for speed. We first calculate the 1D center of each bin in the range $-1$ to $1$. For each of the $n$ annotations of activation, we subtract the center of each bin from the annotation, resulting in a $64$ size vector, which we call $x$. The contribution to the $64$ bins will then be calculated using an element-wise $sigmoid$ on this vector, $sigmoid(\sigma*(x+\frac{\delta}{2}))-sigmoid(\sigma*(x-\frac{\delta}{2}))$. The gradient of $sigmoid$ is largest at zero, for values of $x$ far from 0, the $\frac{\delta}{2}$ term has less effect, and the bin value is close to 0. This function is maximized for values of $x$ close to 0. We repeat this for valence to get two $n\times 64$ matrices for activation and valence.
In the equation, $\sigma$ is a scaling parameter; the larger the value, the more sharp the histogram is, and $\delta$ is the bin size. Since our data is in the range $-1$ to $1$, and we use 64 bins, $\delta=\frac{2}{64}$. We then matrix multiply these two $n\times 64$ matrices by transposing one to get a 2D ($64\times 64$) matrix. We then normalize to get a final $4\times 4$ probability distribution as in [7]. There is a tradeoff where too large of a $\sigma$ may lead to vanishing gradients, but too low may result in undersaturation [23]. As such, we set $\sigma$ relatively small at $8$; lower values did not reduce loss. Future work could investigate the impact of modifying the $\sigma$ parameter. The generation of probabilities in DiffKDE is done in float16555We use float64 during validation and testing for KDE accuracy. as it significantly speeds up calculations.
@@ 3.4 Evaluation Metrics
We first evaluate the ability of the proposed approaches to learn continuous predictions and then the ability of the system to learn distributions. The baseline cannot directly produce continuous ratings, while the proposed approaches can. In order to provide a fair comparison, we generate consensus predictions across all methods in the same manner: we sum along the activation/valence dimensions and then multiply this sum with $[-1,-0.5,0.5,1]$. We use CCC to measure the systems’ ability to predict individual annotators’ labels (note: we cannot evaluate the baseline for this task). Next, we evaluate the consensus predictions by comparing them to the averaged ground truth using CCC. Finally, we measure the differences between the learned and ground truth probability distributions using Total Variation Distance (TVD), and Jensen-Shannon Divergence (JSD) [7] [^9] We use natural logarithm for JSD instead of $\log_{2}$. Test results are reported over five seeds. Significance asserted at a 5% confidence on a paired two-sided t-test.
@@ 4 Results
Table 1: MSP-Improv probability distribution results (*=statistical significant improvement compared to baseline, $\dagger$=statistical significant decline). $\uparrow$ indicates higher is better, $\downarrow$ indicates lower is better. Each metric’s best result is bolded.
@@ 4.3 Cross-Corpus Results
Table 2: Cross-corpus zero-shot Activation, Valence results, *$\dagger$,$\uparrow$,$\downarrow$ as in Table 1. P: MSP-Podcast, I: IEMOCAP, M: MuSE

### v2

@@ 2 Related Work
Instead, we build on the label processing method developed in previous work by Zhang et al. [7], which incorporated inter-annotator variance into machine learning models by creating new ground truth labels that incorporate this knowledge [7]. They upsampled existing annotations by selecting random subsets of annotators for each utterance and took the mean across those annotations. They added random noise to the resulting means, such that $x\ noise\sim U(-\frac{std(x)}{2},\frac{std(x)}{2})$ [^1] We also add $\epsilon=1E^{-12}$ to this value to account for cases where standard deviation is 0., where $x$ is activation or valence, stdstd indicates the standard deviation of the annotator ratings for that utterance, and $U$ is the uniform distribution. Kernel Density Estimation (KDE) via Diffusion was then calculated over the upsampled observations. They divided the KDE output grid into $N$ bins for each dimension and took the mean over the KDE samples inside each bin. They converted this grid to a probability distribution by normalizing over the means. The authors investigated $N=2$ and $N=4$. The KDE step was essential to remove sensitivity to where boundaries were drawn. The authors then trained a model to predict these binned distributions. However, in this approach, the model loses information about individual annotators. Additionally, because the approach is not differentiable, it cannot be included in model training. We present an approach with a differentiable component that permits learning a binned distribution, implemented using sigmoid-based soft operations.
@@ 3.1 Data setup
Dataset Preprocessing We process all datasets in the same way. We use min-max scaling on the annotator and consensus labels for activation and valence to restrict labels to the [1,1][-1,1] range. We then use KDE to generate a 2D ground truth probability distributions as in [7]. We use a KDE grid size of 512 as we assume this will be sufficiently large to ensure the probability is insensitive to the grid boundaries.
@@ 3.2 Model Architecture
The baseline model directly learns the generated KDE distribution (as in [7]), having a final linear layer output of 16 logits for the 4x4 discretized KDE distribution prediction case. The MT model has separate prediction layers for each annotator as in [17]. Each annotator's continuous prediction of activation and valence is made via a linear layer with an output size of 1. We use the one-hot method, previously used for text emotion recognition [20]. We use the same architecture as in the MT case but with only one annotator prediction head. The annotator ID is one-hot encoded and concatenated to the Wav2Vec2 and BERT embeddings on the model input. When training within corpus, we reduce computation cost by making predictions only for annotators in the batch input to the model.
@@ 3.3 Training Tasks
In this section, we define three different training tasks. The Baseline model is trained with the Baseline task. MT and one-hot models are trained by interleaving Tasks 1 and 2 (Task 1+2), defined below. We use stochastic gradient descent with a learning rate of $0.001$, with a learning rate scheduler that adjusts the learning rate by a factor of $0.1$ after five epochs of no reduction in validation metrics. We train models until early stopping triggers with a patience of 10 with a minimum of 30 epochs. Each model trains with a batch size of 32. For all methods we use the relevant task's validation losses.
@@
Task 1 - Annotator Training: We train annotator-specific predictions using the individual annotator ground truth. We use Lin's Concordance Correlation Coefficient (CCC) loss as our loss function since it better models dimensional attributes than other regression losses [21]. The sets actact, and valval contain the ground truth labels from all annotators in a training batch. The sets $m_{act}$ and $m_{val}$ contain the model's estimates of these labels. The loss is $2-CCC(m_{act},act)-CCC(m_{val},val)$.
Task 2 - DiffKDE: We learn the probability distribution using the KDE-generated ground truth labels. The model must produce a probability distribution from the model's activation and valence predictions. However, the KDE method outlined by Zhang et al. in [7] is not immediately usable. KDE via diffusion starts with a histogram [22]. For each annotation we must know if it is in a particular bin to increment the bin's histogram count. This operation is a binary operation and not differentiable. We introduce a differentiable approximation to this problem by instead calculating a confidence value that a given annotation is within a given bin. We modify an existing one-dimensional (1D) soft-histogram [^3] https://discuss.pytorch.org/t/differentiable-torch-histc/25865/4, as below, for the 2D data.
We use 64 bins for DiffKDE [^4] Note: smaller than 512 (used to generate target labels) for speed. We first calculate the 1D center of each bin in the range 1-1 to $1$. For each of the $n$ annotations of activation, we subtract the center of each bin from the annotation, resulting in a $64$ size vector, which we call $x$. The contribution to the $64$ bins will then be calculated using an element-wise sigmoidsigmoid on this vector, $sigmoid(\sigma*(x+\frac{\delta}{2}))-sigmoid(\sigma*(x-\frac{\delta}{2}))$. The gradient of sigmoidsigmoid is largest at zero, for values of $x$ far from 0, the $\frac{\delta}{2}$ term has less effect, and the bin value is close to 0. This function is maximized for values of $x$ close to 0. We repeat this for valence to get two $n\times 64$ matrices for activation and valence.
In the equation, $\sigma$ is a scaling parameter; the larger the value, the more sharp the histogram is, and $\delta$ is the bin size. Since our data is in the range 1-1 to $1$, and we use 64 bins, $\delta=\frac{2}{64}$. We then matrix multiply these two $n\times 64$ matrices by transposing one to get a 2D ($64\times 64$) matrix. We then normalize to get a final $4\times 4$ probability distribution as in [7]. There is a tradeoff where too large of a $\sigma$ may lead to vanishing gradients, but too low may result in undersaturation [23]. As such, we set $\sigma$ relatively small at $8$; lower values did not reduce loss. Future work could investigate the impact of modifying the $\sigma$ parameter. The generation of probabilities in DiffKDE is done in float16555We use float64 during validation and testing for KDE accuracy. as it significantly speeds up calculations.
@@ 3.4 Evaluation Metrics
We first evaluate the ability of the proposed approaches to learn continuous predictions and then the ability of the system to learn distributions. The baseline cannot directly produce continuous ratings, while the proposed approaches can. In order to provide a fair comparison, we generate consensus predictions across all methods in the same manner: we sum along the activation/valence dimensions and then multiply this sum with [1,0.5,0.5,1][-1,-0.5,0.5,1]. We use CCC to measure the systems' ability to predict individual annotators' labels (note: we cannot evaluate the baseline for this task). Next, we evaluate the consensus predictions by comparing them to the averaged ground truth using CCC. Finally, we measure the differences between the learned and ground truth probability distributions using Total Variation Distance (TVD), and Jensen-Shannon Divergence (JSD) [7] [^9] We use natural logarithm for JSD instead of $\log_{2}$. Test results are reported over five seeds. Significance asserted at a 5% confidence on a paired two-sided t-test.
@@ 4 Results
Table 1: MSP-Improv probability distribution results (*=statistical significant improvement compared to baseline, $\dagger$=statistical significant decline). $\uparrow$ indicates higher is better, $\downarrow$ indicates lower is better. Each metric's best result is bolded.
@@ 4.3 Cross-Corpus Results
Table 2: Cross-corpus zero-shot Activation, Valence results, $\dagger$$\uparrow,\downarrow$ as in Table 1. P: MSP-Podcast, I: IEMOCAP, M: MuSE